OpenStudy (anonymous):

what is the simplest form of 16/24 (ratio)

OpenStudy (anonymous):

Divide 4 on the top and bottom, 4 is the largest # they both share

OpenStudy (♪chibiterasu):

A.K.A The GCF.

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

16/4 = 4 24/4=6

OpenStudy (anonymous):

so im guessing 4:6 is the answer

OpenStudy (anonymous):

Oh crap, you can divide by 8 :P

OpenStudy (kendricklamar2014):

\[16/8 = 2\] \[24/8 = 3 \] \[Answer = \left(\begin{matrix}2 \\ 3\end{matrix}\right)\]

OpenStudy (anonymous):

Giving answers? @KendrickLamar2014

OpenStudy (kendricklamar2014):

8 is the G.C.F

OpenStudy (kidrah69):

Worked showed

OpenStudy (kendricklamar2014):

I'm not giving direct answers, I was showing him how to solve the problem using the G.C.F

OpenStudy (♪chibiterasu):

Even if the work is shown it is still considered a direct answer.

OpenStudy (anonymous):

@kidrah69 what is the gcf of 4 and 20

OpenStudy (kidrah69):

You tell me

OpenStudy (anonymous):

idk thats why im asking for help

OpenStudy (kidrah69):

Well how do you find the gcf?

OpenStudy (anonymous):

You should look up how to find the GCF... @filipinochick

OpenStudy (♪chibiterasu):

You're questions are just what he is asking. @kidrah69

OpenStudy (kidrah69):

Nope. If he cant find the gcf or atleast learn it then its not possible to do these problems which is why they should learn that first http://www.mathsisfun.com/greatest-common-factor.html

OpenStudy (anonymous):

im a SHE

OpenStudy (anonymous):

key word CHICK

OpenStudy (anonymous):

LOL ¯\_(ツ)_/¯

OpenStudy (♪chibiterasu):

Well why don't you teach her instead of judging her and linking her somewhere else.

OpenStudy (kidrah69):

Thats the link to help her...

OpenStudy (anonymous):

Lets all be Friends ʕ•ᴥ•ʔ

OpenStudy (♪chibiterasu):

:|

OpenStudy (kidrah69):

Chbi is my friend he just doesnt know who i am :P

OpenStudy (♪chibiterasu):

I am completely aware of who you are.

OpenStudy (anonymous):

Lol ~(˘▾˘~)

OpenStudy (kidrah69):

Cool then :P @filipinochick did u read the link?

OpenStudy (kendricklamar2014):

\[O-O\]

OpenStudy (anonymous):

it didnt help

OpenStudy (anonymous):

... it explains everything

OpenStudy (♪chibiterasu):

See? Linking them somewhere else isn't effective. It's better to explain it yourself.

OpenStudy (anonymous):

i learn best if someone explains it too me i dont do well off of videos

OpenStudy (kidrah69):

There is no video

OpenStudy (anonymous):

\(\large\cal\color{#ff5d00}{The ~website~ explains~ it ~o.o}\)